OpenStudy (anonymous):

What is the quotient of 3.33 × x10^–5 and 0.06?

OpenStudy (aum):

\[\large \frac{3.33 \times 10^{-5}}{0.06} = \frac{3.33 \times 10^{-5}}{6 \times 10^{-2}} = \frac{3.33 }{6 \times 10^{-2} \times 10^{5} } = \frac{0.555 }{10^{3} } = \\ \large \frac{0.555 \times 10}{10^{3}\times 10 } = \frac{5.55}{10^{4} } = 5.55 \times 10^{-4} \]